Solve the system of equations
2x - y + 3z = -2
x + 4y - 2z = 16
5x + y - 1z = 14

Answer by Alan3354(69443)

Solve the system of equations
2x - y + 3z = -2
x + 4y - 2z = 16
5x + y - 1z = 14
================
Multiply Eqn 3 by 2, then subtract from Eqn 2.
-----
Multiply Eqn 3 by 3, then add to Eqn 1.
================
z is eliminated.
---
etc.
Email via the TY note for help or to check your work.